OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 64x^2

OpenStudy (btaylor):

f(x) = (x)(x)(x^2 + 64) f(x) = (x)(x)(x+8i)(x-8i)

OpenStudy (anonymous):

can you show me the steps please?

OpenStudy (ash2326):

We have \[f(x)=x^4+64x^2\] first step is to take x^2 common from both the terms, we get \[f(x)=x^2(x^2+64)\] Do you get this @madixy ?

OpenStudy (anonymous):

yes! :)

OpenStudy (ash2326):

Have you studied complex no.s ?

OpenStudy (anonymous):

yes but I still have trouble with "i"

OpenStudy (btaylor):

basically, think of it as a difference of squares: x^2+64 = x^2-(-8)^2

OpenStudy (btaylor):

^ + -8^2

OpenStudy (ash2326):

we know that \[\sqrt{-1} =i\] I'd rewrite f(x) as \[f(x)=x^2(x^2-(-64))\] do you get this point?

OpenStudy (anonymous):

yes i get it now! thank you!

OpenStudy (ash2326):

could you solve the rest?

OpenStudy (anonymous):

x^2(x + 8i)(x - 8i)

OpenStudy (ash2326):

Good work:D

OpenStudy (anonymous):

thanks again!